6.A 25.0-mL sample of chlorine gas (Cl2), with a temperature of 15 °C, is warmed up to 35 °C. What is its new volume?
Answer #1

6.

From ideal gas equation ,

V1/T1 = V2/T2

(25/(15+273.15)) = (V2/(35+273.15))

v2 = volume of gas at 35 C = 26.735 ml